# Understanding Biases in ChatGPT-based Recommender Systems: Provider Fairness, Temporal Stability, and Recency

### v1

@@ 3. Evaluation of ChatGPT-based RecLLM
Our work focuses on evaluating ChatGPT as a zero-shot recommender in the movie domain, aiming to extend beyond previous research by delving more deeply into biases and potential harms caused by such systems. Given the extensive nature of our experiments, we chose to concentrate on the single domain of movie recommendations. In this context, our evaluation methodology for ChatGPT was comprehensive and multifaceted. We conducted extensive experiments across a multi-dimensional testbed to assess the robustness and adaptability of the system. This evaluation particularly emphasized understanding the role and impact of ChatGPT in three key areas:
  1. (1)
    Goal-oriented prompt,
  2. (2)
    System role,
  3. (3)
    Promoting fairness, as well as
  4. (4)
    Implcit vs. explicit scenario
These dimensions were varied to observe their individual and combined effects on the quality of the recommendations provided. Please note that since the primary objective of this study is centered on (provider) fairness, the emphasis on fairness has been integrated either as a distinct scenario within the goal-oriented prompts or within the system role itself, and the impact is evaluated.
3.1. Goal-oriented prompts
Table 2. Overview of Prompt Scenarios.
@@ 3.1. Goal-oriented prompts
S1 - Simple
Basic recommendations without additional context.
@@
S2 - Genre-focused
Recommendations focusing on genres and themes similar to the user’s past favorites.
@@
S3 - Rating-focused
Incorporation of the user’s explicit ratings into the recommendations.
@@
S4 - Diversify Recommendations
Suggesting lesser-known films to diversify the user’s experience.
@@
S5 - Surprise
Offering unexpected recommendations for exploring new preferences.
@@
S6 - Motivate Reasoning
Providing reasoning for each recommendation to enhance transparency.
@@
S7 - Chain of Thought (COT)
Engaging in a logical reasoning process to arrive at the recommendations.
@@
In particular, scenarios S1 to S3 are focused on core personalization aspects, encompassing basic recommendations without additional context (S1), genre preferences (S2), and the incorporation of explicit user ratings (S3). Meanwhile, Scenarios (S4) and (S5) are designed to enhance the diversity and novelty of the recommendations, with S4 emphasizing lesser-known film recommendation and S5 aiming to broaden the user’s cinematic horizons with unique and unexpected choices. Scenarios S6 and S7 are explanation-motivation oriented, where S6 seeks to provide reasoning for each recommendation to enhance transparency, and S7 involves a logical ste-by-step reasoning process to arrive at the recommendations. These scenarios compose a competitive and insightful set of prompts, some of which such as COT, have been successfully tested in other ML disciplines (Wei et al., 2022; Wang et al., 2022; Chu et al., 2023).
Note. The CF baselines used in this work operate in an implicit setting. For consistency and fairness, all scenarios in the generative part adhered to this setting, thereby not revealing user movie ratings to the LLM in the prompts. Scenario ‘Rating-focused’ S3 is the only exception, included solely for completeness. This is discussed in Section 3.5.
3.2. Repeated Experiment for the Stability of the Analysis
The experiments were conducted at a temperature setting of 0.0, aiming for predictable responses. However, we observed variations in the recommendations, possibly due to factors like trending data, recent updates in the database, or changes in GPT’s behavior. This suggests that the recommendation system may interpret the same input differently in each iteration, independent of the temperature setting.
To assess the robustness and reliability of our system, we repeatedly conducted recommendation queries for each user, five times each, to examine if outputs vary despite consistent input.
We further adopted a “bootstrapping sampling strategy” during the evaluation stage, as detailed in Section 4.3. This method is particularly effective in revealing the reliability of the results, even when dealing with datasets that are small or may not perfectly represent the entire population. This thorough evaluation approach, therefore, enhances our confidence in the system’s performance across various potential real-world scenarios.
3.3. Understanding ChatGPT’s “System” Role
In this study, we sought to understand the impact of different system roles assigned to ChatGPT on recommendation outcomes. We focused on assessing whether responses significantly differ when specific “system” instructions are attached, and how these variations influence the functionality of a recommender system. Specifically, we aimed to determine the effectiveness of embedding fairness directly into the system’s operations versus incorporating it in the prompt structure. Table 3 summarizes the roles assigned to ChatGPT during our experiments:
@@ 3.3. Understanding ChatGPT’s “System” Role
R0 - No Role
Direct user-driven prompts without system context.
R1 - System Role as Recommender
System-centric role focusing purely on user information.
R2 - System Role as Fair Recommender
System role with an explicit fairness objective.
The roles explored provide insight into how ChatGPT’s framing, in terms of its responsibilities and objectives, shapes its behavior in making recommendations. This analysis is crucial for comprehending the subtleties of human-AI interaction within diverse contexts and for distinct purposes, especially in how the system interprets and adheres to varying directives.
3.4. Promoting Fairness
@@
3.5. Explicit vs. Implicit Scenario
This test evaluated whether displaying the movies a user has rated, as indicated in {user_movies_string}, enhances the recommender system’s quality, based on the premise that including such information could improve recommendations. Scenarios S1 to S6, except S3, were conducted implicitly, without user ratings in the prompts, mirroring the implicit mode of the CF baseline experiments. Uniquely, S3 was examined in both explicit and implicit contexts. In the explicit context, user ratings were integrated into the prompts, but were absent in the implicit context. This explicit approach was specific to S3. An example of {user_movies_string} and the presentation of user movie lists and ratings in the explicit context is as follows.
@@
The primary hypothesis under investigation is whether the inclusion of explicit ratings enables the system to tailor recommendation more closely to the user’s demonstrated preferences, thereby enhancing the personalization of the recommendation process.
@@ 4. Experimental Setup
This section outlines experiments conducted to assess various reproducibility aspects, including the obtained results and our observations.
4.1. Datasets and splitting
In this study, we utilized the MovieLens-Latest-Small dataset provided by the GroupLens group, [^1] https://grouplens.org/datasets/movielens/ which contains the most recent versions of movies. It is important to note that we intentionally focused our research within the domain of movies. This decision was made due to the potential varying capabilities of GPT-based models across different domains. Additionally, conducting extensive experiments with GPT models is time and resource-intensive, which is why we opted to work with this specific dataset. We acknowledge that there is potential for future exploration in other domains, and we leave this avenue open for future research. Detailed dataset parameters, including $\frac{R}{U}$ (ratings per user) and $\frac{R}{I}$ (ratings per item), resulting in a higher density $\frac{R}{U\times I}$, are presented in Table 4 for reference. Our setting involves a train-validation-test split for the data with ratios 80%, 10%, and 10%, respectively.
4.2. Evaluation Metrics
Our study employs a diverse range of metrics to assess the performance of our recommendation system. These metrics are categorized into two primary classes: Accuracy, Beyond-accuracy and Provider Fairness.
Accuracy Metrics. These metrics measure the quality of the top-$k$ ranking in the recommendation list.
  • NDCG@k. A measure of ranking quality, particularly the placement of relevant items.
@@ 4.2. Evaluation Metrics
Beyond-Accuracy and Provider Fairness Metrics. Beyond conventional accuracy metrics, these metrics evaluate the diversity, and fairness of the recommendations w.r.t provider perspective. They assess how well the system introduces new and varied content and how equitably it treats different item providers.
  • Gini Index: Measures the inequality in the distribution of item recommendations. A lower Gini Index indicates a more equitable distribution among items. The Gini Index is beneficial for identifying monopolization of recommendations by a few items, but it may not always capture the full picture of item diversity.
  • HHI (Herfindahl-Hirschman Index): Assesses the concentration of recommendations among items. A lower HHI value suggests a healthier, more diverse recommendation landscape. This index is particularly useful for detecting dominance by a small number of items, though it might overlook the distribution nuances among less recommended items.
  • Entropy: Reflects the diversity and unpredictability of recommendations. Higher entropy values indicate greater diversity. Entropy is advantageous for understanding the overall unpredictability of recommendations but may not directly indicate the fairness of item exposure.
Each of these metrics has its own advantages and limitations, making them collectively essential for a well-rounded evaluation. They help in understanding not just the effectiveness of the recommendation system in terms of user satisfaction (accuracy) but also its robustness in promoting diverse and fair content distribution.
4.3. Bootstrapping Sampling Strategy
In our evaluation stage, we employed a bootstrapping sampling strategy, a statistical technique where 1000 samples were generated by repeatedly resampling our dataset with replacement. This method was chosen to enhance the robustness and reliability of our analysis, particularly in measuring the performance metrics of our recommendation system, such as NDCG, Recall, and Precision.
Through this approach, we computed means for each metric across all bootstrap samples, then calculated their averages and 95% confidence intervals. This strategy is implemented for understanding the variability and confidence in our metrics, ensuring our evaluation reflects the system’s likely real-world performance. It is especially beneficial in situations with small or non-representative samples, allowing for more reliable population inferences.
4.4. CF Baselines.

### v3

@@ 2.2. Leveraging Pre-trained LMs and Prompting for Recommender Systems
In recent surveys (Deldjoo et al., 2024a, b), the wide area of generative models has been extensively explained. These surveys highlight the potential of generative models in transforming traditional RS by leveraging their capabilities in understanding and generating complex data distributions. For instance, they have shown how generative models can effectively handle multimodal data (text, images, and videos), thereby enhancing the richness and accuracy of recommendations. These surveys also shed light on the emergent reasoning abilities of LLMs, such as in-context learning, which allows models to adapt to new tasks with minimal additional training. Moreover, they discuss the application of retrieval-augmented generation (RAG) techniques, which combine information retrieval with generative modeling to produce contextually relevant recommendations. Overall, these insights underscore the significant advancements and ongoing challenges in utilizing generative models for RS.
@@ 3. Evaluation of ChatGPT-based RecLLM
We designed two experiments to evaluate the performance and bias of ChatGPT-based RecLLMs.
  • Experiment 1. Examining prompt design strategies in classical top-$k$ recommendation (cf. Section 3.1), and
  • Experiment 2. Studying generalization of findings across datasets, and learning strategies (zero vs. few-shot ICL) in sequential recommendation task (cf. Section 3.2)
Table 2 provides a summary of these experiments.
Table 2. Summary of Experiments
Exps. Aim Approach Evaluation Metrics
Exp 1 (i) Investigate prompt design strategies in classical top-K recommendations (ii) Analyze stability of results over different RecLLM runs, (iii) Look at new content-related metrics (genre diversity, temporal freshness) Explore prompt formulations (accuracy, diversity, fairness), system role directive, specific information in prompt in the movie domain Accuracy (top-K precision and recall), beyond-accuracy metrics (diversity, novelty), fairness metrics, genre diversity, temporal freshness
Exp 2 (i) Assess generalization of fairness across datasets and beyond zero-shot learning to ICL, (ii) Look at the impact of user profile construction and demographic information revealing Use movie and music dataset, evaluate few-shot and zero-shot ICL in sequential recommendation Accuracy, provider fairness, catalogue coverage
3.1. Experiment 1: Examining Prompt Design Strategies in Classical Top-K Recommendations
In the first experiment, we investigate prompt design strategies in a classical top-$K$ recommendation setting. We use an LLM (ChatGPT) as a zero-shot recommender, querying it with various prompt formulations to obtain recommendations. Our goal is to understand how much the recommendation outcome (performance) can be influenced by prompts carrying different objectives (e.g., relevance, increasing diversity, or motivating reasoning). ur goal is to determine if a one-to-one mapping exists between the desired natural language query (e.g., enhance diversity) and the actual diversity of the model recommendations.
To achieve the goals of this paper, we examine the relevance and item fairness (including diversity) of the recommendations. Additionally, we explore new aspects of recommendation, such as content diversity and freshness, including genre diversity and temporal freshness.
Table 3. Overview of prompt scenarios designed for the experiment 1.
@@ 3.1. Experiment 1: Examining Prompt Design Strategies in Classical Top-K Recommendations
S1 - Simple Basic recommendations without additional context.
@@
S2 - Genre-focused Recommendations focusing on genres and themes similar to the user’s past favorites.
@@
S3 - Rating-focused Incorporation of the user’s explicit ratings into the recommendations.
@@
S4 - Diversify Recommendations Suggesting lesser-known films to diversify the user’s experience.
@@
S5 - Surprise Offering unexpected recommendations for exploring new preferences.
@@
S6 - Motivate Reasoning Providing reasoning for each recommendation to enhance transparency.
@@
S7 - Chain of Thought (COT) Engaging in a logical reasoning process to arrive at the recommendations.
@@
3.1.1. Goal-oriented prompts
In particular, scenarios S1 to S3 are focused on core personalization aspects, encompassing basic recommendations without additional context (S1), genre preferences (S2), and the incorporation of explicit user ratings (S3). Meanwhile, Scenarios (S4) and (S5) are designed to enhance the diversity and novelty of the recommendations, with S4 emphasizing lesser-known film recommendation and S5 aiming to broaden the user’s cinematic horizons with unique and unexpected choices. Scenarios S6 and S7 are explanation-motivation oriented, where S6 seeks to provide reasoning for each recommendation to enhance transparency, and S7 involves a logical ste-by-step reasoning process to arrive at the recommendations. These scenarios compose a competitive and insightful set of prompts, some of which such as COT, have been successfully tested in other ML disciplines (Wei et al., 2022b; Wang et al., 2022b; Chu et al., 2023). [^1] Note. The CF baselines used in this work operate in an implicit setting. For consistency and fairness, all scenarios in the generative part adhered to this setting, thereby not revealing user movie ratings to the LLM in the prompts. Scenario ‘Rating-focused’ S3 is the only exception, included solely for completeness. This is discussed in Section 3.1.5.
3.1.2. Repeated Experiment for the Stability of the Analysis
The experiments were conducted at a temperature setting of 0.0, aiming for predictable responses. However, we observed variations in the recommendations, possibly due to factors like trending data, recent updates in the database, or changes in the behavior of ChatGPT. This suggests that the recommendation system may interpret the same input differently in each iteration, independent of the temperature setting.
To assess the robustness and reliability of our system, we repeatedly conducted recommendation queries for each user, five times each, to examine if outputs vary despite consistent input. We further adopted a bootstrapping sampling strategy during the evaluation stage, as detailed in Section 4.1.3. This method is particularly effective in revealing the reliability of the results, even when dealing with datasets that are small or may not perfectly represent the entire population. This thorough evaluation approach, therefore, enhances our confidence in the performance of the system across various potential real-world scenarios.
3.1.3. Understanding the impact of “System” Role in ChatGPT
In this study, we sought to understand the impact of different system roles assigned to ChatGPT on recommendation outcomes. We focused on assessing whether responses significantly differ when specific “system” instructions are attached, and how these variations influence the functionality of a recommender system. Specifically, we aimed to determine the effectiveness of embedding fairness directly into the system role versus incorporating it in the prompt itself. Table 4 summarizes the roles assigned to ChatGPT during our experiments:
@@ 3.1.3. Understanding the impact of “System” Role in ChatGPT
R0 - No Role Direct user-driven prompts without system context.
R1 - System Role as Recommender System-centric role focusing purely on user information.
R2 - System Role as Fair Recommender System role with an explicit fairness objective.
3.1.4. Fairness Emphasis
@@
3.1.5. Explicit vs. Implicit Scenario
This test evaluated whether displaying the movies a user has rated, as indicated in {user_movies_string}, enhances the quality of RS, based on the premise that including such information could improve recommendations. Scenarios S1 to S6, except S3, were conducted implicitly, without user ratings in the prompts, mirroring the implicit mode of the CF baseline experiments. Uniquely, S3 was examined in both explicit and implicit contexts. In the explicit context, user ratings were integrated into the prompts, but were absent in the implicit context. This explicit approach was specific to S3. An example of {user_movies_string} and the presentation of user movie lists and ratings in the explicit context is as follows.
@@
The primary hypothesis under investigation is whether the inclusion of explicit ratings enables the system to tailor recommendations more closely to the user’s demonstrated preferences, thereby enhancing the personalization of the recommendation process.
Figure 2. Sequential in-context learning for various scenarios explored in Experiment 2 of this research.
3.2. Experiment 2: Sequential In-Context Learning
The task of recommending suitable movies/music to users can be modeled as a sequential ranking problem where the historical interactions of a user are utilized to predict future preferences. In this context, the goal is to rank candidate items $C=\{i_{j}\}_{j=1}^{m}$ such that the most relevant items for the user appear at the top. Given a user’s interaction history $H=\{i_{1},i_{2},\ldots,i_{n}\}$, to leverage it for generating recommendations, we opt to divide it into two main parts, as displayed in Figure 2:
  1. (1)
    Context part. This part is sampled from the entire interaction history of the user to provide general context about the user’s interests, passion, and preference (toward music and movies). We use two strategies for sampling:
    • Frequent/Highest-rated: Selects the most enjoyed items by the user, based on the ratings provided (for movies) or frequency of consumption (for music);
    • Recency-Focused: Utilizes a weighted average method to emphasize recently consumed items. This strategy selects items that have not only been liked by the user but also have been recently listened to or watched, ensuring relevance and timeliness in the recommendations.
  2. (2)
    Example part. This part is taken from recent interactions to guide specific next recommendations. It directly determines the next items to be consumed in sequential tasks and is used in ICL-$1$, and ICL-$2$, which denote the few-shot learning scenarios, using one or two examples.
Based on this division, we explore different prompting strategies, as outlined in the following. We note that Experiment 2 has been inspired by Hou et al. (2024) and extends it by exploring the above profile investigation, and evaluation going beyond pure accuracy.
Note. In the prompting scenarios outlined below, we optionally incorporate “user demographic information” (such as gender, or age category) into the prompts to examine how revealing this information might influence model biases and stereotypes. Specifically, where we have {is demographic} information, it is used in statements like ‘‘The user is female’’, ‘‘The user is young’’. Details of the results of comparisons with and without demographic information on fairness are provided in Table 10.
Sequential Zero-Shot Prompting
This method only uses the context part of the user interaction history
Example:
The user {is female, and} has watched the following movies:
- "From Russia with Love (1963)" in genre(s) Action with rating 5
- "Star Trek IV: The Voyage Home (1986)" in genre(s) Action|Adventure|Sci-Fi with rating 4
- "Planet of the Apes (1968)" in genre(s) Action|Sci-Fi with rating 3
- "Star Wars: Episode I - The Phantom Menace (1999)" in genre(s) Action|Adventure|Fantasy|Sci-Fi with rating 4
- "Final Conflict, The (a.k.a. Omen III: The Final Conflict) (1981)" in genre(s) Horror with rating 3
This selection reflects the user’s movie preferences.
What would be the top-1 suitable next recommendation?
as stated earlier, {is female} is essentially denoted with {is demographic} below, which is tested in two scenarios: included and omitted.
Sequential In-Context Learning (ICL-1)
This approach takes the contest part of the preivous part, but taked the xample part of the interaction and from that take the last interatcion to guide the model how the recommendation. The examples are derived from recent interactions.
Example:
The user {is demographic, and} has watched the following movies:
- "From Russia with Love (1963)" in genre(s) Action with rating 5
- "Star Trek IV: The Voyage Home (1986)" in genre(s) Action|Adventure|Sci-Fi with rating 3
- "Planet of the Apes (1968)" in genre(s) Action|Sci-Fi with rating 3
- "Star Wars: Episode I - The Phantom Menace (1999)" in genre(s) Action|Adventure|Fantasy|Sci-Fi with rating 4
- "Final Conflict, The (a.k.a. Omen III: The Final Conflict) (1981)" in genre(s) Horror with rating 3
This selection reflects the user’s movie preferences.
Given the user has recently watched the following movies in order:
1. "Face/Off (1997)" in genre(s) Action|Sci-Fi|Thriller with rating 4
2. "Bringing Out the Dead (1999)" in genre(s) Drama|Horror with rating 5
3. "Sixth Sense, The (1999)" in genre(s) Thriller with rating 5
4. "Austin Powers: The Spy Who Shagged Me (1999)" in genre(s) Comedy with rating 4
5. "Arlington Road (1999)" in genre(s) Thriller with rating 4
You should recommend:
Recommendation 1: "Bowfinger (1999)" in genre(s) Comedy with rating 3
What would be the top-1 suitable next recommendation after the above movies?
Sequential In-Context Learning (ICL-2)
Similar to ICL-1 but with a bigger selection of examples equal to 2. The examples are chosen to provide more specific guidance on generating recommendations based on recent user interactions.
Example:
The user {is demographic, and} has watched the following movies:
- "From Russia with Love (1963)" in genre(s) Action with rating 5
- "Star Trek IV: The Voyage Home (1986)" in genre(s) Action|Adventure|Sci-Fi with rating 3
- "Planet of the Apes (1968)" in genre(s) Action|Sci-Fi with rating 3
- "Star Wars: Episode I - The Phantom Menace (1999)" in genre(s) Action|Adventure|Fantasy|Sci-Fi with rating 4
- "Final Conflict, The (a.k.a. Omen III: The Final Conflict) (1981)" in genre(s) Horror with rating 3
This selection reflects the user’s movie preferences.
Given the user has recently watched the following movies in order:
1. "Strange Days (1995)" in genre(s) Action|Crime|Sci-Fi with rating 3
2. "Face/Off (1997)" in genre(s) Action|Sci-Fi|Thriller with rating 4
3. "Bringing Out the Dead (1999)" in genre(s) Drama|Horror with rating 5
4. "Sixth Sense, The (1999)" in genre(s) Thriller with rating 5
5. "Austin Powers: The Spy Who Shagged Me (1999)" in genre(s) Comedy with rating 4
You should recommend:
Recommendation 1: "Bowfinger (1999)" in genre(s) Comedy with rating 3
Recommendation 2: "Arlington Road (1999)" in genre(s) Thriller with rating 4
What would be the top-1 suitable next recommendation after the above movies?
Recency-Focused Zero-shot and ICL-1 and ICL-2
We replicate all previous prompting scenarios but modify the “context part” to include the most recent interactions. This approach indicates to the ChatGPT recommender system that the best recommendations are based on the user’s recent consumption.
The next sections outlines experiments conducted to assess various reproducibility aspects, including the obtained results and our observations.
@@ 4. Experimental Setup
In this section, we detail the experimental setup for two experiments, including the datasets, evaluation metrics, bootstrapping methods, baselines employed, and hyperparameter tuning.
4.1. Experiment 1. Classical Top-$k$ Recommendation
4.1.1. Datasets and Setup
Experiment 1 used the MovieLens-Latest-Small dataset from the GroupLens group. [^2] https://grouplens.org/datasets/movielens/ This dataset comprises the latest movie entries and was used in its entirety to carry out experiments in part 1 of our study. The experimental design required conducting API calls for each user, resulting in a total of 610 API calls (equal to the number of users) per individual recommendation setting. Given the scope—610 users, 7 different prompts or prompt scenarios (cf. Section 3.1.1), 2 fairness statements in the prompt (cf. Section 3.1.4), and 2 system roles (cf. Section 3.1.3)—we generated thousands of scenarios or API calls. Due to the computational demands, we limited our experiments in part 1 to a single dataset, focused on the classical top-$k$ recommendation task. The detailed statistics of the dataset are provided in Table 5, where we implemented a data split of 80% training, 10% validation, and 10% testing. The validation dataset is used solely for CF baseline hyperparameter tuning, while both RecLLMs and the CF baseline utilize the same training and testing datasets to conduct recommendations.
4.1.2. Evaluation Metrics
The first experiment employs a diverse range of metrics to assess the performance of different RS. These metrics are categorized into two primary classes: accuracy metrics, and provider fairness/diversity. Additionally, we incorporate new content-centric metrics such as temporal freshness and genre dominance/tendancy to explore insights into potential biases and tendencies of RecLMMs compared to mainstream CF models.
Accuracy Metrics.
These conventional metrics measure the quality of the top-$k$ recommendation or the personalization of recommendations
  • NDCG@k: Measures the ranking quality by evaluating the placement of relevant items within the top-$k$ positions.
@@ 4.1.2. Evaluation Metrics
Provider Fairness/Diversity. Beyond conventional accuracy metrics, these metrics evaluate the fairness of the recommendations from the provider perspective. They assess how well the system introduces new and varied content and how equitably it treats different item providers.
  • Gini Index (Deldjoo et al., 2021b): Measures the inequality in the distribution of item recommendations. A lower Gini Index indicates a more equitable distribution among items. The Gini index is calculated as follows:
    $\text{Gini}=\frac{\sum_{i=1}^{n}(2i-n-1)x_{i}}{n\sum_{i=1}^{n}x_{i}}$
    where $x_{i}$ represents the number of times item $i$ is recommended, sorted in ascending order, and $n$ is the total number of items. The Gini index becomes 0 when all items are recommended an equal number of times, indicating perfect equality. Mathematically, if all $x_{i}$ are equal, then:
    $x_{i}=\frac{\sum_{i=1}^{n}x_{i}}{n}\quad\forall i\implies\text{Gini}=0$
    The Gini index becomes 1 when one item is recommended exclusively, indicating maximum inequality. In our work, we interpret the Gini index to measure item fairness, where lower Gini values indicate higher (better) item fairness.
  • HHI (Herfindahl-Hirschman Index) (of Justice, 2018; Wang et al., 2022a): This metric Assesses the concentration of recommendations among items. It is calculated as follows:
    $\text{HHI}=\sum_{i=1}^{n}\left(\frac{x_{i}}{\sum_{j=1}^{n}x_{j}}\right)^{2}$
    The HHI value ranges from $1/n$ to 1, where $n$ is the total number of items. The HHI is $1/n$ when all items are recommended equally, indicating maximum fairness, and it is 1 when one item is recommended exclusively, indicating no fairness. Here, lower values of the HHI are considered indicative of better item fairness.
  • Entropy: Reflects the diversity and unpredictability of recommendations. Higher entropy values indicate greater diversity. The Entropy is calculated as follows:
    $\text{Entropy}=-\sum_{i=1}^{n}p_{i}\log(p_{i})$
    where $p_{i}=\frac{x_{i}}{\sum_{j=1}^{n}x_{j}}$ and $p_{i}>0$. The entropy value ranges from 0 to $\log(n)$, where 0 indicates no diversity (one item recommended exclusively) and $\log(n)$ indicates maximum diversity (all items recommended equally). Higher values of Entropy are considered indicative of better item fairness.
While the above metrics are used to measure item fairness based on item exposure, they also serve as indicators of how recommendations produce diverse outcomes.
4.1.3. Bootstrapping Sampling Strategy
In experiment 1, for the evaluation, we additionally employ a bootstrapping sampling strategy, a statistical technique where 1000 samples were generated by repeatedly resampling our dataset with replacement (Deldjoo, 2023). This method was chosen to enhance the robustness and reliability of our analysis, particularly in measuring the performance metrics of our recommendation system, such as NDCG, Recall, and Precision.
Through this approach, we computed means for each metric across all bootstrap samples, then calculated their averages and 95% confidence intervals. This strategy is implemented for understanding the variability and confidence in our metrics, to ensure our evaluation reflects the likeness of the system performance on real-world performance. This is especially beneficial in situations with small or non-representative samples, allowing for more reliable population inferences.
4.1.4. CF Baselines.